OpenStudy (anonymous):

\[{ 1 \over R} = {1 \over R1} + {1 \over R2} \]

OpenStudy (anonymous):

Solve the equation for the indicated variable.

OpenStudy (anonymous):

guess it depends on the indicated variable.

OpenStudy (anonymous):

\[{1 \over R}\] ? lol

OpenStudy (anonymous):

are you solving for \[R\] \[R_1\] or \[R_2\]?

OpenStudy (anonymous):

no. i'm sorry. it's R1

OpenStudy (anonymous):

whew. ok

OpenStudy (anonymous):

haha. work with me satellite. :p

OpenStudy (anonymous):

subtract \[\frac{1}{R_2}\] from both sides, take the reciprocal

OpenStudy (anonymous):

\[\frac{1}{R}-\frac{1}{R_2}=\frac{1}{R_1}\]

OpenStudy (anonymous):

\[{1 \over R} - {1 \over R2} = {1 \over R1} ? \]

OpenStudy (anonymous):

\[\frac{R_2-R}{RR_2}=\frac{1}{R_1}\]

OpenStudy (anonymous):

haha. same thing! this is good!

OpenStudy (anonymous):

now flip everything to get \[\frac{RR_2}{R_2-R}=R_1\]

OpenStudy (anonymous):

where did you come up with the last one?

OpenStudy (anonymous):

the one before the last... not flipping but the previous one

OpenStudy (anonymous):

oh i had 1/R and wanted R. just flip it

OpenStudy (anonymous):

oh previous one. ok hold on

OpenStudy (anonymous):

\[\frac{1}{a}-\frac{1}{b}=\frac{b}{ab}-\frac{a}{ab}\] yes? common denominator is ab and i have to build up the fractions

OpenStudy (anonymous):

then you subtract to get \[\frac{b-a}{ab}\]

OpenStudy (anonymous):

clear?

OpenStudy (anonymous):

no. give me a second. i'm going to write it...

OpenStudy (anonymous):

ok if not clear let me know. think of \[\frac{1}{5}-\frac{1}{3}\]

OpenStudy (anonymous):

forgot to close your tag

OpenStudy (anonymous):

this: \[\frac{1}{a}-\frac{1}{b}=\frac{b}{ab}-\frac{a}{ab} \] is important for me to remember. i am taking note of it. i think it's just that i need to understand this before being able to do these kind of problems. does that sound right from what you are seeing with me?

OpenStudy (anonymous):

but i understood

OpenStudy (anonymous):

yes, but don't get married to that formula. to add and subtract fraction it is always \[\frac{a}{bb}\pm\frac{c}{d}=\frac{ad\pm bc}{bd}\]

OpenStudy (anonymous):

thanks for confusing me more :p

OpenStudy (anonymous):

just in this case the numerators happened to be 1!

OpenStudy (anonymous):

that should really not confuse you much. how do you add \[\frac{2}{3}+\frac{5}{7}\]? it is just \[\frac{2\times 7 + 3\times 5}{3\times 7}\]

OpenStudy (anonymous):

i would rather find the common denominator for the fraction with numbers lol

OpenStudy (anonymous):

but you are showing me a more advanced way. i believe i must learn this way to be succesful

OpenStudy (anonymous):

in your case it was \[\frac{1}{R}-\frac{1}{R_2}=\frac{1\times R_2-1\times R}{R\times R_2}\]

OpenStudy (anonymous):

so the 1's were eliminated.... then we are left with variables

OpenStudy (anonymous):

wtf. that's the easiest way to add fractions that i have seen.

OpenStudy (anonymous):

thank you again! so the: \[\frac{a}{bb}\pm\frac{c}{d}=\frac{ad\pm bc}{bd} \] applies to all fractions?